# Directed Exploration in Reinforcement Learning from Linear Temporal Logic

### v1

@@ Appendix D Additional Tabular Results
Optimal policies for the reach-avoidance tasks from Figure 4 (first row) would also solve a shortest-path reaching problem in the underlying MDP. In other words, a linear path to the goal does not violate the constraints. This Section evaluates a variation of the original task, by reshaping the safe area of the reach-avoid task into a U-shape. In this case, the shortest path to the goal violates the avoidance constraints. Fortunately, changing the shape of the avoidance zone only results in a local change to the labeling function $\mathcal{F}$. DRL2, as well as other baselines, can handle arbitrary relabeling functions; thus these methods can adapt to arbitrary environment layouts.
Results are reported in 7. for all task variants (easy, medium and hard). In general, as the path length has now increased, performance globally decreases accordingly. Nevertheless, the overall relative performance of each methods is consistent with the original version of the task. We note that LCER [45] now outperforms the count-based baseline, as it can effectively ignore previous visits to the zone to avoid. In the extreme case in which the length of the U-maze is very large, but the starting position is very close to the goal if the avoidance constraint is ignored, we expect LCER to perform even better.
@@ Appendix E Comparison to LTL-guided Counterfactual Experience Replay (LCER)
LTL-guided Counterfactual Experience Replay [44] takes an orthogonal approach to the problem of exploration when learning from LTL formulas and represents a very strong baseline in specific environments. In particular, LCER does not leverage the semantics of the LDBA, however it generates synthetic experience by fully controlling its dynamics. Transitions $((s,b),a,(s^{\prime},b^{\prime}))$ with $\{(s,b),(s^{\prime},b^{\prime})\}\subseteq\mathcal{S}^{\times}$ and $a\in\mathcal{A}^{\times}$ are relabeled by uniformly resampling the initial LDBA state and simulating the LDBA transition caused by action $a$, resulting in a new transition $((s,\hat{b}),a,(s^{\prime},\hat{b}^{\prime}))$.
Generated transitions respect the transition kernel of the product MDP and are thus valid training samples. However, the method does not guarantee that relabeled product states $(s,\hat{b})$ are actually feasible, as the set of reachable product states might only consist of a manifold of the full Cartesian product of MDP and LDBA state spaces. For instance, let us consider the illustrative task $T_{0}$ as displayed in Figure 2: at any step, if the agent is in the red field, the LDBA would necessarily be in its sink state and in no other state. Any relabeling of this particular state would thus not be relevant. Furthermore, as a form of state relabeling, LCER cannot generate on-policy data, as the on-policy action sampled in $(s,\hat{b})$ might be different than the original action sampled in $(s,b)$ [34].
@@ Appendix F Kernel Estimation Ablation
  • a transition kernel obtained using value iteration (VI), and that therefore represents the optimal policy under assumption of full controllability of an high level MDP. It represents a naive implementation of the reward shaping introduced in Icarte et al. [22]. This kernel has the lowest entropy possible, as value iteration outputs a deterministic policy.
@@ Appendix G Prior Strength Ablation
Appendix H Extended Related Works
Logic for Task Specification
The intersection of reinforcement learning and logic languages has flourished around the topic of Reward Machines [7, 43, 22]. While Reward Machines can handle regular expressions, LTL is designed to address a superset, namely $\omega$-expressions. As a consequence, LTL gains the ability to reason about infinite sequences and tasks including liveness and stability.
For this reason, several works have investigated LTL for task specification. Early attempts proposed heuristic reward shaping schemes [28], which gradually evolved towards dynamic or eventual rewarding and discounting schemes [17, 18, 9, 8, 5, 45]. As for the optimization of such schemes, while linear programming is sufficient for known dynamics and finite action spaces, Q-learning approaches have been adopted when a model of the environment is not available [17, 4, 5]. A fundamental restriction of this approaches was lifted by adopting actor-critic [18] or policy gradient [45] methods suited for continuous action spaces. Nevertheless, in this case, empirical demonstration of LTL-guided RL have remained mostly constrained to low-dimensional setting.
Logic-driven Exploration in Reinforcement Learning
A significant effort has been targeted at addressing the sparsity arising from proposed rewarding schemes [17, 45].
Perhaps closest to the method we propose, several works investigate reward shaping for reinforcement learning from logic specification. In the context of reward machines, Camacho et al. [6] consider potential functions $\Phi(s,b)$ of both the MDP and automaton states; among those, they investigate counting the minimum number of steps to reach an accepting automaton state, which is equivalent to computing values under a specific automaton transition kernel.
Another work [22] explores an extension, which is also fundamentally related to our method: the authors investigate value iteration on an MDP synthetised from the automaton to compute a potential function for reward shaping, thus leveraging the semantics of the task. While motivated by a similar intuition to ours, a naive application of this method would not be viable in an eventually discounted setting, which involves $\omega$-regular problems, as we report in Appendix F. By adopting an LDBA instead of a DFA, and more importantly by integrating an eventual discounting scheme, our method is able to avoid myopic behavior while handling the full complexity of LTL. Futhermore, DRL2 does not require assumptions on controllability by defining a Markov reward process instead. Values estimated in the MRP are thus softer than the optimal ones in a corresponding MDP, and we find them to be more informative for exploration. We validate this discussion empirically, by evaluating optimal transition kernels under eventual discounting in Appendix F. By updating its distribution over MRP transition kernels, it is moreover capable of providing an adaptive and informative reward signal, rather than a static one derived from a hard assumption.
@@ Appendix I Evaluation with $\epsilon$-transitions
This section evaluates application in presence of $\epsilon$-transitions. $\epsilon$-transitions occur in a subset of automata, in particular for stability specification; our method, as well as baselines considered, are compatible with theis existence. Thus, we now evaluate the task $FGa$ in reach-avoid tabular settings. This setup is similar to that of the corresponding experiment in recent work [45]. Overall, as the structure of the LDBA is minimal, we observe that all considered methods perform reasonably well in Figure 11. In particular, DRL2 and the count-based baseline slightly improve over the No-exploration baseline. LCER [45] works particularly well, as the task does not involve several substeps, but presents irrecoverable failures when "jumping" at the wrong time. We refer to Appendix E for a further comparison between DRL2 and LCER.
@@
Figure 11: Evaluation of DRL2 and baselines on a the task $FGa$ in presence of jump transitions.
Appendix J Implementation Details
This section presents a thorough description of methods, tasks and hyperparameters. For the sake of reproducibility, we provide our full codebase 333sites.google.com/view/drl-2.
J.1 Environments and Tasks
@@
J.1.2 Continuous Environments
Section 4 also provides an evaluation on continuous, high-dimensional environments. Since evaluation of LTL-guided RL has traditionally been mostly restricted to simple tabular [22] or low-dimensional [18, 45] settings, we repurpose high-dimensional environments from safe and goal-conditioned reinforcement learning research.
@@
Experiments reported in the first row of Figure 5 were developed on top of the common Fetch robotic benchmark [10], and are also available as part of our codebase. We evaluate two tasks, in which the agent controls the end effector position of a 7DoF robotic arm with 4-dimensional actions over episodes of 50 steps. FetchAvoid is reminiscent of the reach-avoidance task in tabular settings, and involves fully stretching out the arm while avoiding lateral movements (i.e., not entering red zones as shown in Figure 5). In this case, observations are 10-dimensional. In FetchAlign the agent has to interact with three cubes on one side of the table, and has to align them horizontally at the center of the table. In this case, observations include information on the cubes, and are 45-dimensional. The two tasks are specified, respectively, with the formulas $Fa\&G\neg e$ (reach the end of the table, and avoid lateral movements), and $F(a\&XF(b\&XFc))$ (position the first, second and then third block).
@@
Doggo is a 12DoF quadruped adapted from the most challenging tasks in SafetyGym [35]. It is tasked with navigating a flat plane; observations are 66-dimensional, actions are 12-dimensional and the length of each episode is 500 steps. Similarly to other reach-avoidance tasks, DoggoAvoid involves navigation to a distant goal (shown in green in Figure 5), in a straight line, while completely avoiding detours. On the other hand, DoggoLinear involves navigating through a sequence of 2 circular zones. These two tasks are encoded through the following two specifications: $Fa\&G\neg e$ and $F(a\&XFb)$. Episode length is set to 500.
@@
HalfCheetah [42] involves controlling a 6DoF robot in a vertical 2D plane and is a standard environments in the deep reinforcement learning literature. We define two tasks in this underlying environments. In CheetahSequential the agent is tasked with eventually reaching, in succession, 4 consecutive zones to its right, as encoded by the formula $F(a\&XF(b\&XF(c\&XFd)))$. Once the last zone is reached, the task is solved and the agent can stop. In CheetahFrontflip the agent is informed by the formula $GF(a\&XF(b\&XF(c\&XFd)))$, where each variable evaluates to true for a given range of angles of the main body of the robot, respectively when the Cheetah is in its default position, standing on its front legs, upside down, and standing on its back legs. As a result, the formula is satisfied by an infinite sequence of frontflips. Notably, this task can not be satisfied by finite trajectories.
J.2 Methods and Algorithms
The core of our method computes an intrinsic reward and is therefore applicable on top of arbitrary reinforcement learning algorithms. Therefore, we will first discuss the implementation of the method itself, as well as relevant baselines, and then details involving RL algorithms.
@@
The main hyperparameter introduced by DRL2 is $\alpha$, which controls the strength of its prior as described in Section 3.2. It is set to $10^{3}$ in tabular settings, and to $10^{5}$ in continuous settings, where increased stability was found to be beneficial. The remaining hyperparameters for reward shaping are shared with the count-based baseline and are, respectively, the frequency of updates to the potential function (set to 2000 environment steps), and a scaling coefficient to the intrinsic reward, which was tuned individually for each method in a grid of $[0.1,1.0,10.0]$. We found a coefficient of 0.1 to be optimal across all tasks for both methods in tabular settings, while continuous settings benefit from a stronger signal and a coefficient of $10.0$. The two remaining baselines (relying on LCER and no exploration techniques) introduce no additional hyperparameters on top of the learning algorithm.
@@
The intrinsic reward is added to the extrinsic one and optimized under eventual discounting through one of two algorithms, depending on the setting. In practice, we rely on tabular Q-learning [48] with an $\epsilon$-greedy policy ($\epsilon$=0.1) and Soft Actor Critic with automatic entropy tuning [14]. In the tabular case, strict adoption of eventual discounting [45] results in a largely uniform policy, as in most states no action would reduce the likelihood of task satisfaction. As a consequence, a prohibitive number of steps would be required during evaluation for computing cumulative returns. For this reason, our practical implementation of tabular Q-learning does not apply eventual discounting, although this would be possible given a much larger computational budget. Our SAC implementation is partially based on that provided by Huang et al. [19]. Hyperparameters for each algorithm were tuned to perform best when no exploration bonus is being used, and are reported in Tables 1. They are kept fixed across tasks.
J.3 Metrics
@@
Unless specified, the metric used to measure the methods’ performance is cumulative return under eventual discounting, that is $R=\sum_{i=0}^{N}\gamma^{i}$, where $\gamma=0.99$ and $N$ is the number of visits to an accepting LDBA state within an episode. In the limit of $N\to\infty$, this metric is a proxy for the likelihood of task satisfaction [45]. All curves report mean performances estimated across 10 seeds, and shaded areas represent $95\%$ simple bootstrap confidence intervals. They undergo average smoothing with a kernel size of 10.
J.4 Tools
@@
Our codebase [^4] Available at sites.google.com/view/drl-2. mostly relies on numpy [16] for numeric computation and torch [30] for its autograd functionality. Furthermore, we partially automate the synthesis of LDBAs from LTL formulas through rabinizer [25].
J.5 Computational Costs
@@
All methods have comparable runtimes within their setting (tabular or continuous). Each experimental run required 8 cores of a modern CPU (Intel i7 12th Gen CPU or equivalent). Each run in tabular and continuous settings required on average  30 minutes and  150 minutes, respectively.

### v2

@@ Appendix D Additional Tabular Results
UMaze
Optimal policies for the reach-avoidance tasks from Figure 4 (first row) would also solve a shortest-path reaching problem in the underlying MDP. In other words, a linear path to the goal does not violate the constraints. We now evaluate a variation of the original task, by reshaping the safe area of the reach-avoid task into a U-shape. In this case, the shortest path to the goal violates the avoidance constraints. Fortunately, changing the shape of the avoidance zone only results in a local change to the labeling function $\mathcal{F}$. DRL2, as well as other baselines, can handle arbitrary relabeling functions; thus these methods can adapt to arbitrary environment layouts.
Results are reported in 7. for all task variants (easy, medium and hard). In general, as the path length has now increased, performance globally decreases accordingly. Nevertheless, the overall relative performance of each methods is consistent with the original version of the task. We note that LCER (Voloshin et al., 2023) now outperforms the count-based baseline, as it can effectively ignore previous visits to the zone to avoid. In the extreme case in which the length of the U-maze is very large, but the starting position is very close to the goal if the avoidance constraint is ignored, we expect LCER to perform even better.
@@
Sequential maze
We additionally evaluate an alternative layout of the sequential task, which considers a greater number of APs, and in which the sequence of zones to visit leads to the exit of a maze with several intersections (see Figure 8). We scale the difficulty by increasing the areas of each zone. As the reward signal synthesized by DRL2 mostly depends on the structure of the LDBA, we observe consistent results with the sequential tasks in Figure 4.
  DRL2  LCER    Count-based    No exploration
Figure 8: Alternative evaluation of a sequential task in the shape of a maze. Results are consistent with those in Figure 4.
@@ Appendix E Comparison to LTL-guided Counterfactual Experience Replay (LCER)
LTL-guided Counterfactual Experience Replay (Voloshin et al., 2022) takes an orthogonal approach to the problem of exploration when learning from LTL formulas and represents a very strong baseline in specific environments. In particular, LCER does not leverage the semantics of the LDBA, however it generates synthetic experience by fully controlling its dynamics. Transitions $((s,b),a,(s^{\prime},b^{\prime}))$ with $\{(s,b),(s^{\prime},b^{\prime})\}\subseteq\mathcal{S}^{\times}$ and $a\in\mathcal{A}^{\times}$ are relabeled by uniformly resampling the initial LDBA state and simulating the LDBA transition caused by action $a$, resulting in a new transition $((s,\hat{b}),a,(s^{\prime},\hat{b}^{\prime}))$.
Generated transitions respect the transition kernel of the product MDP and are thus valid training samples. However, the method does not guarantee that relabeled product states $(s,\hat{b})$ are actually feasible, as the set of reachable product states might only consist of a manifold of the full Cartesian product of MDP and LDBA state spaces. For instance, let us consider the illustrative task $T_{0}$ as displayed in Figure 2: at any step, if the agent is in the red field, the LDBA would necessarily be in its sink state and in no other state. Any relabeling of this particular state would thus not be relevant. Furthermore, as a form of state relabeling, LCER cannot generate on-policy data, as the on-policy action sampled in $(s,\hat{b})$ might be different than the original action sampled in $(s,b)$ (Rauber et al., 2019).
@@ Appendix F Kernel Estimation Ablation
  • a transition kernel obtained using value iteration (VI), and that therefore represents the optimal policy under assumption of full controllability of an high level MDP. It represents a naive implementation of the reward shaping introduced in Icarte et al. (2022). This kernel has the lowest entropy possible, as value iteration outputs a deterministic policy.
@@ Appendix G Prior Strength Ablation
Appendix H Product MDP Ablation: States and Rewards
The product MDP formulation that our method builds upon (Voloshin et al., 2023) can be used to formally connect eventually discounted returns and likelihoods of formula satisfaction, but strictly requires a binary $1/0$ reward signal. Moreover, the agent state space is defined as the product between LDBA and MDP states. In this section, we empirically evaluate trends that arise if these two properties are violated.
  DRL2  No exploration    LDBA-only    Reward-shift
Figure 12: Evaluation of additional baselines in easy tabular tasks: LDBA-only trains a policy directly on the LDBA/MRP state, and Reward-shift shifts rewards to {$/1\}$ 1 }.
States
In principle, the agent needs access to both high-level LDBA states (tracking formula satisfaction) and low-level MDP states, tracking fine-grained details. While this is necessary to provably recover an optimal policy for arbitrary MDPs and formulas, it also causes a blow-up of the state space, which makes exploration more difficult. As rewards are only defined as a function of the LDBA states, one might be tempted to directly learn a policy (and value function) from LDBA states alone $\pi:\mathcal{B}\to\mathcal{A^{\times}}$. This decision has the large advantage of operating over a much smaller state space, and the disadvantage of losing information about low-level MDP information. This is evident in the empirical evaluation we present for tabular tasks in Figure 12 (see curves for LDBA-only. Across all tasks, the optimal value function is not invariant with respect to the MDP state; thus, value estimates are inaccurate, and the policy does not solve any of the tasks for the entirety of training.
Rewards
The eventual discounting framework manages to retrieve a proxy to an infinite-horizon objective by adopting a discount factor of exactly $1$ for non-accepting transitions. As a consequence, the reward function for such transitions needs to be exactly zero. If this assumption is violated, the value function of a policy which does not visit an accepting transition infinitely often will diverge to (possibly negative) infinity. For this reason, from a formal perspective, the binary reward signal is crucial. However, in practical setting, one can apply an affine transformation to the reward signal, shifting it as $R^{\prime}(b)=2R(b)-1\in\{-1,1\}$, thus punishing each timestep that does not transition to an accepting state. We evaluate this heuristic solution in tabular settings and observe interesting results in Figure 12. In simpler tasks (e.g., reach-avoid), this solution performs very well. This can be traced back to the fact that the Q-values are initialized to zero: while under $R$ they are not updated until an accepting state is visited by chance, in this case they will decrease as soon as the corresponding action is executed. This encourages the agent to prioritize yet unattempted actions, driving efficient exploration. This effect leads this baseline to match the performance of DRL2 in simpler tasks, but is less effective in more complex tasks (e.g., circular), in which struggles to learn and is outperformed by the existing No-exploration baseline.
@@ Appendix I Evaluation with $\epsilon$-transitions
This section evaluates application in presence of $\epsilon$-transitions. $\epsilon$-transitions occur in a subset of automata, in particular for stability specification; our method, as well as baselines considered, are compatible with theis existence. Thus, we now evaluate the task $FGa$ in reach-avoid tabular settings. This setup is similar to that of the corresponding experiment in recent work (Voloshin et al., 2023). Overall, as the structure of the LDBA is minimal, we observe that all considered methods perform reasonably well in Figure 13. In particular, DRL2 and the count-based baseline slightly improve over the No-exploration baseline. LCER (Voloshin et al., 2023) works particularly well, as the task does not involve several substeps, but presents irrecoverable failures when "jumping" at the wrong time. We refer to Appendix E for a further comparison between DRL2 and LCER.
@@
Figure 13: Evaluation of DRL2 and baselines on the task $FGa$ in presence of jump transitions.
  DRL2  LCER    Count-based    No exploration
Figure 14: Evaluation of DRL2 and baselines on the task $Fa\&G\neg b$ with stochastic initial states.
Appendix J Evaluation under Stochasticity
Stochastic initial state
Experiments in Figure 4 initialize the agent in a fixed state. Our evaluation does not however depend on this assumption. In order to verify this, we repeat the evaluation for the reach-avoid task, and sample the initial state uniformly from a rectangle in the leftmost part of the "corridor". In Figure 14, results are consistent with those in the original setting, with all methods being slightly more sample-efficient as they can be initialized in a more favorable position in some episodes.
Sticky actions
We additionally repeat the tabular evaluation while injecting a difference source of stochasticity: sticky actions. Inspired by evaluation protocols on ATARI, with probability $p=0.1$ at each step, we force the environment to ignore the agent’s action and simply repeat the action from the previous time-step. This induces both stochasticity, and a slight partial observability, as the transitions now depend on previous actions as well. We report results in Figure 15. We observe that the overall trends from the original, deterministic variants are are largely reproduced, with an expected overall slight drop in performance. This suggests that the evaluated methods are at least in part robust to stochasticity.
  DRL2  LCER    Count-based    No exploration
Figure 15: Evaluation of DRL2 and baselines on medium tabular tasks from Figure 4 with sticky actions.
Appendix K Evaluation under On-policy Learning
DRL2 is in principle agnostic to the underlying RL algorithm, as it relies simply on reward shaping. We further validate this property empirically, by re-evaluating tabular problems of medium difficulty from Figure 4, and replacing Q-learning for SARSA Sutton & Barto (2018). In Figure 16, we observe that the overall sample efficiency decreases for all method, which is consistent with the on-policy nature of SARSA. Nevertheless, the relative performance of all methods remains consistent with the results obtained with Q-learning.
  DRL2  LCER    Count-based    No exploration
Figure 16: Repetition of tabular experiments while replacing Q-learning with SARSA.
  DRL2  LCER    Count-based    No exploration
Figure 17: Evaluation of DRL2 and baselines on the task $Fa\&G\neg b$ in terms of violations of the avoidance constraind $G\neg b$.
Appendix L Additional Metrics
Contraint violations
While the standard evaluation in Figure 4 reports returns under eventual discounting, this section additionally reports the ratio of training episode that violate the avoidance constraints in the task $Fa\&G\neg b$. In Figure 17 we observe that, as expected, this metric is largely negatively correlated with evaluation returns: in order to maximize the likelihood of formula satisfaction, constraint violations need to be minimized. We additionally note one interesting trend: the count-based baseline quickly compensates the high rate of violations for a randomly initialized policy and quickly minimizes the violations. However, it requires more training steps to also reliably reach the specified zone.
Formula satisfaction in finite time
Our empirical evaluation mainly relies on eventually discounted returns as a proxy for formula satisfaction. However, on some of the evaluation tasks, we can evaluate an additional metric. Namely, on sequential tasks without avoidance constraints, it is sufficient to visit an acceptance state once to satisfy the formula. While estimating the likelihood of visitation over infinite horizons remain intractable, we can evaluate the likelihood of a policy to visit an accepting state (and therefore satify the formula) within the episode length. We report an empirical estimate of this quantity in Figure 18, obtained by computing bootstrap confidence intervals of a satisfaction indicator across seeds. We observe that variance over seeds is limited: therefore, the plot closely resembles a clipped version of the metrics reported in Figure 4.
  DRL2  LCER    Count-based    No exploration
Figure 18: Evaluation of the likelihood of formula satisfaction in finite time for sequential tasks from Figure 4. DRL2 satisfies the formulas from the earliest episodes, while its eventually discounted return grows during the initial stage of training. Intuitively, this means that the agent reaches an accepting state earlier and earlier.
Appendix M Implementation Details
This section presents a thorough description of methods, tasks and hyperparameters. For the sake of reproducibility, we provide our full codebase 666github.com/marbaga/drl2.
M.1 Environments and Tasks
@@
M.1.2 Continuous Environments
Section 4 also provides an evaluation on continuous, high-dimensional environments. Since evaluation of LTL-guided reinforcement learning has traditionally been mostly restricted to simple tabular (Icarte et al., 2022) or low-dimensional (Hasanbeig et al., 2020, Voloshin et al., 2023) settings, we repurpose high-dimensional environments from safe and goal-conditioned reinforcement learning research.
@@
Experiments reported in the first row of Figure 5 were developed on top of the common Fetch robotic benchmark (de Lazcano et al., 2023), and are also available as part of our codebase. We evaluate two tasks, in which the agent controls the end effector position of a 7DoF robotic arm with 4-dimensional actions over episodes of 50 steps. FetchAvoid is reminiscent of the reach-avoidance task in tabular settings, and involves fully stretching out the arm while avoiding lateral movements (i.e., not entering red zones as shown in Figure 5). In this case, observations are 10-dimensional. In FetchAlign the agent has to interact with three cubes on one side of the table, and has to align them horizontally at the center of the table. In this case, observations include information on the cubes, and are 45-dimensional. The two tasks are specified, respectively, with the formulas $Fa\&G\neg e$ (reach the end of the table, and avoid lateral movements), and $F(a\&XF(b\&XFc))$ (position the first, second and then third block).
@@
Doggo is a 12DoF quadruped adapted from the most challenging tasks in SafetyGym (Ray et al., 2019). It is tasked with navigating a flat plane; observations are 66-dimensional, actions are 12-dimensional and the length of each episode is 500 steps. Similarly to other reach-avoidance tasks, DoggoAvoid involves navigation to a distant goal (shown in green in Figure 5), in a straight line, while completely avoiding detours. On the other hand, DoggoLinear involves navigating through a sequence of 2 circular zones. These two tasks are encoded through the following two specifications: $Fa\&G\neg e$ and $F(a\&XFb)$. Episode length is set to 500.
@@
HalfCheetah (Towers et al., 2023) involves controlling a 6DoF robot in a vertical 2D plane and is a standard environments in the deep reinforcement learning literature. We define two tasks in this underlying environments. In CheetahSequential the agent is tasked with eventually reaching, in succession, 4 consecutive zones to its right, as encoded by the formula $F(a\&XF(b\&XF(c\&XFd)))$. Once the last zone is reached, the task is solved and the agent can stop. In CheetahFrontflip the agent is informed by the formula $GF(a\&XF(b\&XF(c\&XFd)))$, where each variable evaluates to true for a given range of angles of the main body of the robot, respectively when the Cheetah is in its default position, standing on its front legs, upside down, and standing on its back legs. As a result, the formula is satisfied by an infinite sequence of frontflips. Notably, this task can not be satisfied by finite trajectories.
M.2 Methods and Algorithms
The core of our method computes an intrinsic reward and is therefore applicable on top of arbitrary reinforcement learning algorithms. Therefore, we will first discuss the implementation of the method itself, as well as relevant baselines, and then details involving reinforcement learning algorithms.
@@
The main hyperparameter introduced by DRL2 is $\alpha$, which controls the strength of its prior as described in Section 3.2. It is set to $10^{3}$ in tabular settings, and to $10^{5}$ in continuous settings, where increased stability was found to be beneficial. The remaining hyperparameters for reward shaping are shared with the count-based baseline and are, respectively, the frequency of updates to the potential function (set to 2000 environment steps), and a scaling coefficient to the intrinsic reward, which was tuned individually for each method in a grid of $[0.1,1.0,10.0]$. We found a coefficient of 0.1 to be optimal across all tasks for both methods in tabular settings, while continuous settings benefit from a stronger signal and a coefficient of $10.0$. Intuitively, higher coefficients can lead to better value propagation during the exploration phase at the cost of stability in deep learning settings; however, we found their impact to be limited (especially in tabular settings), such that no task-specific tuning was required. The two remaining baselines (relying on LCER and no exploration techniques) introduce no additional hyperparameters on top of the learning algorithm.
@@
The intrinsic reward is added to the extrinsic one and optimized under eventual discounting through one of two algorithms, depending on the setting. In practice, we rely on tabular Q-learning (Watkins & Dayan, 1992) with an $\epsilon$-greedy policy ($\epsilon$=0.1) and Soft Actor Critic with automatic entropy tuning (Haarnoja et al., 2018b). In the tabular case, strict adoption of eventual discounting (Voloshin et al., 2023) results in a largely uniform policy, as in most states no action would reduce the likelihood of task satisfaction. As a consequence, a prohibitive number of steps would be required during evaluation for computing cumulative returns. For this reason, our practical implementation of tabular Q-learning does not apply eventual discounting, although this would be possible given a much larger computational budget. Our SAC implementation is partially based on that provided by Huang et al. (2022). Hyperparameters for each algorithm were tuned to perform best when no exploration bonus is being used, and are reported in Tables 1. They are kept fixed across tasks.
M.3 Metrics
@@
Unless specified, the metric used to measure the methods’ performance is cumulative return under eventual discounting, that is $R=\sum_{i=0}^{N}\gamma^{i}$, where $\gamma=0.99$ and $N$ is the number of visits to an accepting LDBA state within an episode. In the limit of $N\to\infty$, this metric is a proxy for the likelihood of task satisfaction (Voloshin et al., 2023). All curves report mean performances estimated across 10 seeds, and shaded areas represent $95\%$ simple bootstrap confidence intervals. They undergo average smoothing with a kernel size of 10.
M.4 Tools
@@
Our codebase [^7] Available at github.com/marbaga/drl2. mostly relies on numpy (Harris et al., 2020) for numeric computation and torch (Paszke et al., 2017) for its autograd functionality. Furthermore, we partially automate the synthesis of LDBAs from LTL formulas through rabinizer (Křetínskỳ et al., 2018).
M.5 Computational Costs
In principle, the distillation of the intrinsic reward according to DRL2 depends on the size of the LDBA size. As we perform value estimation in the MRP built from the LDBA by computing the closed form solution of the Bellman equation, the complexity of this operation is cubic in the LDBA size. This remains tractable for reasonably sized LDBAs with $<100$ states, and the computational cost is in practice negligible for the LDBAs in our empirical evaluation. For very large LDBAs, sampling-based alternatives for approximate value estimations should be considered instead.
@@
When considering the training times of the underlying RL agent, all methods have comparable runtimes within their setting (tabular or continuous). Each experimental run required 8 cores of a modern CPU (Intel i7 12th Gen CPU or equivalent). Each run in tabular and continuous settings required on average $\sim$30 minutes and $\sim$150 minutes, respectively.